# Very Basic Set theory

### v6

@@ Abstract.
The naive conception of set is described as ”Any predicate has an extension”. Then Russell’s paradox is used to prove the inconsistency of naive set theory. Russell’s paradox reads ”No set can contain all and only those sets which do not contain themselves”. We call a set which contains itself an ‘individual’. In this article we show Russell’s paradox is logically equivalent to ‘for every set $s$, an individual is contained in $s$ or a set that is not an individual is not contained in $s$’. Very basic set theory is a set theory with individuals that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus.
@@ Introduction
In (Oliver and Smiley [2006] pp.145-146) Alex Oliver and Timothy Smiley explain
individuals are not needed therefore are prohibited in set theory.
@@
Thus ‘$x$ $\varepsilon$ $x$ or not’ is definite; the possibility ”$x$ $\varepsilon$ $x$” is not excluded by our axioms.
Therefore Zermelo’s 1908 axiomatization permitted sets which contain themselves.
Sets are denoted by $s$, $t$, $u$, …. Symbol $\in$ is the ‘membership relation’ symbol. Then $s\in t$ means $s$ is a ‘constituent’ (also called a member) of set $t$. The negation of $s\in t$ is denoted by $s\notin t$. The membership relation is definite, that means
@@
Then for every set $s$ it is the case $s\in s$ or $s\notin s$, but not both. We define
  • a set is an ‘individual’ if the only constituent of that set is that set itself.
In Section 1 and Section 2 we describe individuals and sets, in Section 3 we prove logical formulas preserve definiteness. Then Russell’s paradox is proved and axioms specification, equality, pairs, regularity and union are formulated. Some important theorems are proved, and in Section 10 natural numbers are defined.
1. What is an individual?
In set theory $s\in s$ is possible thus individuals are possible. But a set that contains itself, an individual, is problematic (Boolos [1971] p.217, pp.219-220)
@@ 1. What is an individual?
However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. A set does not ‘contain’ something, the constituents of a set exist.
Georg Cantor defined a set as (Boolos [1971] p.215)
  • ”a ‘many, which can be thought of as one, i.e. a totality of definite
    elements that can be combined into a whole by a law’.”
Thus a set is something which can be thought of as one, therefore a set is a ‘mental picture’ in your mind. That is in line with (Wittgenstein [1999] p.39) proposition
  • ”2.1 We make to ourselves pictures of facts.”
Therefore we prefer to write ‘constituent of a set’, not ‘contained in a set’.
Set theory is a language that follows rules of logic to communicate the mental picture of a set with the physical world. Then someone reading, seeing or listening that language forms a mental picture of a set. That picture agrees or disagrees with the picture in your mind. (Wittgenstein [1999] p.43) proposition
  • ”2.222 In the agreement or disagreement of its sense with reality, its truth
    or falsity consists.”
An individual is what is called in (Wittgenstein [1999]) ”a simple sign” (propositions
3.2, 3.201, 3.202, 3.26). In the Introduction (p.12) Bertrand Russell explains
”If an atomic fact is analyzed as fully as possible (theoretical, not practical
possibility is meant) the constituents finally reached may be called simples
or objects.”
”(…) the naming of simples is shown to be what is logically first in logic.”
Therefore in Very basic set theory the naming of individuals is logically first.
2. What is a set?
At an early age you develop the mental picture of ‘what is a set’. You belong to a family or a clan, to the inhabitants of a village or a particular region. Then experience shows there are things that constitute a set. The members of a set are the constituents of that set. Therefore Very basic set theory describes
a world of individuals,
with a definite membership relation between sets,
the basic properties of a set are the constituents of that set and the sets that set is a member of,
the only constituent of an individual is that individual itself.
Experience shows a set ‘is equal to’ or ‘is not equal to’ another set. Sets with different constituents are different sets, and different sets have different constituents.
Set theory follows rules of logic. Logic treats every possibility thus logic treats the definiteness of the membership relation. In the next section we prove the building up of logical formulas preserves definiteness.
@@ 3. Logic and definiteness
A ‘formula’ is a description of properties of sets. Thus $s\in t$ is a formula. Formulas are denoted by $\Phi$, $\Psi$, $\ldots$ and are built up from formulas by means of negation $\neg\Phi$ (not $\Phi$) and conjunction $\Phi\wedge\Psi$ ($\Phi$ and $\Psi$). Negation takes precedence over conjunction. The parenthesis (((( and )))) preclude ambiguity and improve legibility. Then $s\notin t$ is short for the formula $\neg(s\in t)$.
If $\Phi$ is a formula then $(\Phi)$, $\forall s(\Phi)$ and $\exists s(\Phi)$ are formulas. $\forall s(\Phi)$ means ‘for every set it is the case $\Phi$’. The negation of $\forall s(\neg\Phi)$ is denoted by $\exists s(\Phi)$ thus $\exists s(\Phi)$ means ‘exists a set such that $\Phi$’. The parenthesis following $\forall s$ and $\exists s$ are mandatory. However, without ambiguity $\forall s\exists t(s\in t)$ is short for the formula $\forall s(\exists t(s\in t))$.
@@
$\Phi\vee\Psi$ ($\Phi$ or $\Psi$ or both) is short for the formula ¬($\neg(\neg\Phi\wedge\neg\Psi)$,
$\Phi\rightarrow\Psi$ ($\Phi$ implies $\Psi$) is short for the formula ($(\neg\Phi\vee\Psi)$,
@@
There are many rewrite rules: ¬($\neg(\neg\Phi)$ is logically equivalent to $\Phi$, $\neg(\Phi\leftrightarrow\Psi)$ is logically equivalent to $(\Phi\leftrightarrow\neg\Psi)$ and $((\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi))$ is logically equivalent to $((\Phi\wedge\Psi)\wedge(\neg\Phi\vee\neg\Psi))$. More rewrite rules in (Mendelson [2015] p.15).
We define
  • $\Phi$ is definite’ if $\neg(\Phi\wedge\neg\Phi)$ thus it is not the case that both $\Phi$ and $\neg\Phi$.
For every $s$ and for every $t$ it is not the case that both $s\in t$ and $\neg(s\in t)$ thus $s\in t$ is definite We prove: The building up of logical formulas preserves definiteness.
If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise
$(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ or $(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $\exists s(\Phi\wedge\neg\Phi)$,
If $\Phi$ is definite then it is not the case that both $\Phi$ and $\neg\Phi$ thus
it is not the case that both ¬($\neg(\neg\Phi)$ and $\neg\Phi$ therefore $\neg\Phi$ is definite,
$(\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi)$ is logically equivalent to $((\Phi\wedge\neg\Phi)\wedge\Psi)\vee(\Phi\wedge(\Psi\wedge\neg\Psi))$ thus if $\Phi$ and $\Psi$ are definite then $(\Phi\wedge\Psi)$ is definite,
The connectives $\vee$, $\rightarrow$ and $\leftrightarrow$ are expressed in terms of $\neg$ and $\wedge$ thus if $\Phi$ and $\Psi$ are definite then $\Phi\vee\Psi$, $\Phi\rightarrow\Psi$ and $\Phi\leftrightarrow\Psi$ are definite.
Therefore the building up of logical formulas preserves definiteness. $\Box$
4. Russell’s paradox
We prove Russell’s paradox, that is formula $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$
”No set can contain all and only those sets which do not contain themselves”.
Assume exists set $s$ such that for every set $u$ $(u\in s\leftrightarrow u\notin u)$.
The formula applies to every $u$ thus we can choose $s$ therefore
exists $s$ such that $(s\in s\leftrightarrow s\notin s)$.
Formula $(s\in s\leftrightarrow s\notin s)$ is short for $(s\in s\vee s\in s)\wedge(s\notin s\vee s\notin s)$ thus
$(s\in s\wedge s\notin s)$.
Formula $s\in s$ is definite thus it is not the case that $(s\in s\wedge s\notin s)$.
Therefore set $s$ does not exists.
Formula $\exists s\forall u(u\in s\leftrightarrow u\notin u)$ is definite therefore $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$. $\Box$
Then Russell’s paradox $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$ is logically equivalent to
$\forall s\exists u(u\in s\leftrightarrow u\in u)$ thus $\forall s\exists u((u\in s\wedge u\in u)\vee(u\notin s\wedge u\notin u))$.
Therefore the logic of definiteness proves that
for every set $s$, an individual is a constituent of $s$ or a set that is not an individual
is not a constituent of $s$.
5. Axiom ‘specification’
  • if $u$ is free in $\Phi$ and $\Phi$ is definite then
We prove: If an individual is a constituent of $s$ then exists $v$ whose constituents
are all and only those individuals that are constituent of $s$.
Formula $(u\in s\wedge u\in u)$ is definite and is not the conjunction of mutually exclusive properties. Apply ‘specification’ to set $s$ to find set $v$
The formula applies to every $u$ thus we can choose $v$
$\exists v(v\in v\leftrightarrow(v\in s\wedge v\in v))$ therefore
Then $\exists v(v\notin v\vee v\in s)$. Thus $v$ is not an individual, or $v$ is an individual and a constituent of $s$. A set of all individuals is not excluded. $\Box$
We prove: The set of all sets is nonexisting.
If $\exists u(u\in s\wedge u\notin u)$ then formula $(u\in s\wedge u\notin u)$ is definite and is not the conjuction of mutually exclusive properties. Apply ‘specification’ to set $s$ to find set $v$
The formula applies to every $u$ thus we can choose $v$
Then $\exists v(v\notin v\wedge v\notin s)$. Thus $v$ is not an individual and not a constituent of $s$. Therefore the set of all sets is nonexisting. $\Box$
6. Axioms for ‘equality’
@@ 7. Axiom ‘pairs’
Set $v$ is denoted by $\{s,t\}$ thus $\{s,t\}=\{t,s\}$. The ‘singleton’ $\{s\}$ of $s$ is $\{s,s\}$.
@@
However, if $s\neq t$ and $(s\in t\wedge t\in s)$ then $s\notin s$ and $t\notin t$. Then exists set $\{s,t\}$ thus there is the ‘vicious membership circle’ $(s\in t\wedge t\in s)$. That is contradictory to ‘the naming of individuals is logically first’.
8. Axiom ‘regularity’
The naming of individuals is logically first. Thus every set has a constituent, and vicious membership circles do not exist. Therefore we postulate ‘regularity’.
  • (regularity) $\forall s\exists v(v\in s\wedge\forall u((u\in s\wedge u\in v)\rightarrow u\in u))$.
If $s$ is an individual or a set of individuals then $v$ is an individual. Otherwise exists $v$ and $v$ is not an individual and $s$ and $v$ only have individuals in common.
Then it is not the case that exists $s$ such that
$\forall v(v\in s\rightarrow\exists u((u\in s\wedge u\in v)\wedge u\notin u))$.
Therefore any sequence of sets $s$, $t$, $u$, $v$, … , finite (a vicious membership circle) or not finite, such that each term is a member of the previous one
$t\in s$, $u\in t$, $v\in u$ is contradictory to ‘the naming of individuals is logically first’. Then $\forall s\forall t(s\neq t\rightarrow(s\notin t\vee t\notin s))$.
@@ 9. Axiom ‘union’
We prove: Any sequence of different sets $s$, $\bigcup s$, $\bigcup(\bigcup s)$, … is finite.
Then $s\neq\bigcup s\rightarrow\exists t(t\in s\wedge t\notin t)$ otherwise $s=\bigcup s$.
@@
Therefore the sequence terminates thus exists $w$ belonging to the sequence and $\bigcup w=w$. If $z=s\cup\bigcup s\cup\bigcup(\bigcup s)\cup\ldots$ $w$ then $z$ is a set and $\bigcup z=z$. Set $z$ is called ‘the transitive closure’ of $s$. For every set the transitive closure exists. $\Box$
The definition of equality and axiom pairs are only meaningful if exist sets the naming of which are logically first. Thus with individuals and axiom regularity. And regularity is used with axiom union to prove that for every set the transitive closure exists. The proof needs no other axioms and no ‘ set of all natural numbers’.

### v7

@@ Abstract.
The naive conception of set is described as ”Any predicate has an extension”. Then Russell’s paradox is used to prove the inconsistency of naive set theory. If the only constituent of a set is that set itself we call that set an ‘individual’. We show Russell’s paradox permits sets that are individuals thus that is not an inconsistency. Zermelo’s axiomatization also permits individuals. Very basic set theory is a set theory with individuals and has a philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus.
@@ Introduction
Georg Cantor defined a set as (Boolos [1971] p.215)
  • ”a ‘many, which can be thought of as one, i.e. a totality of definite
    elements that can be combined into a whole by a law’.”
That means a set is ‘determined’ by its elements. Sets are denoted by $s$, $t$, $u$, …. Symbol $\in$ is the ‘membership relation’ symbol. Then $s\in t$ means $s$ is a ‘constituent’ (also called ‘an element’ or ‘a member’) of the totality $t$. The negation of $s\in t$ is denoted by $s\notin t$. Set $s$ is an ‘individual’ if the only constituent of $s$ is $s$ itself. A set with only $s$ and $t$ as constituents is denoted by $\{s,t\}$ (Zermelo [1908] p.263). Often the terminology is: ‘set $\{s,t\}$ contains only $s$ and $t$’. But the totality $\{s,t\}$ is ‘thought of as one’ thus a set is a mental picture in your mind. Therefore we prefer to write ‘constituent of a set’, not ‘is contained in a set’.
In (Oliver and Smiley [2006] $\S$5.1) Alex Oliver and Timothy Smiley explain why
”numerous textbooks” prohibit individuals: individuals are not needed for
mathematical purposes.
@@
Thus ‘$x$ $\varepsilon$ $x$ or not’ is definite; the possibility ”$x$ $\varepsilon$ xx” is not excluded by our axioms.
Therefore Zermelo’s 1908 axiomatization permits sets ‘containing’ themselves.
The membership relation is definite, that means
@@
Then for every set $s$ it is the case $s\in s$ or $s\notin s$, but not both. Therefore $s\in s$ is possible. But a set that ‘contains’ itself is considered problematic (Boolos [1971] p.217, pp.219-220)
@@
However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. A set does not ‘contain’ something, the constituents of a set can be thought of as one totality.
Russell’s paradox reads (Boolos [1971] p.217)
”No set can contain all and only those sets which do not contain themselves.”
The negation of Russell’s paradox reads
  • for every $s$ exists a set $u$ such that (($u\in s$ and $u\in u$) or ($u\notin s$ and $u\notin u$)).
If we choose for set $u$ set $s$ then the formula reads ‘for every $s$ ($s\in s$ or $s\notin s$)’. It is the case that ‘for every $s$ ($s\in s$)’ is definite. Therefore Russell’s paradox permits sets ‘containing’ themselves.
In $\S$1 we describe sets and individuals. Atomic facts and elementary propositions are defined which gives the foundation of set theory in the Tractatus (Wittgenstein [1999]). In $\S$2 we prove the formulas of set theory preserve definiteness. In $\S$3 we define equality for sets. Then axioms specification, pairs and union are formulated. We prove regularity and some other important theorems. In $\S$7 natural numbers are defined.
1. What are individuals and sets?
At an early age you develop the mental picture of ‘what is a set’. You belong to a family or a clan, to the inhabitants of a village or a particular region. Then experience shows there are things that constitute a set. The members of a set are the constituents of that set.
Set theory is a language that follows rules of logic to communicate the mental picture of a set with the physical world. Then someone reading, seeing or listening that language forms a mental picture of a set. Set theory is made up of ‘formulas’. A formula is a proposition or predicate about the facts of sets. For every set $s$ and for every set $t$ the fact is the definiteness of $s\in t$. A formula which is not complex or not a combination of other formulas is an ‘atomic fact’. Formula $s\in t$ is an atomic fact, formula $t\in s$ is an atomic fact. Formula $s\notin t$ is the negation of $s\in t$ thus $s\notin t$ is a complex formula therefore $s\notin t$ is not an atomic fact. Atomic facts are what makes formulas true or false.
Definition.
Set $s$ is an individual if the only constituent of set $s$ is set $s$ itself.
A formula presupposes propositions or predicates while propositions or predicates presupposes atomic facts. Then if an atomic fact is analyzed as fully as possible the constituents finally reached are individuals. Individuals cannot be analyzed further. In this way the naming of individuals is what is logically first in logic (Wittgenstein [1999] prop. 3.26) and Bertrand Russell (Introduction p.12).
If $s$ is a set then $s$ is not a proposition or predicate. If $s$ is an individual then $s\in s$ therefore $s$ is the ’propositional connexion’ (Wittgenstein [1999] prop. 4.221), (Anscombe [1963] p.29-31). Thus individuals are ‘elementary propositions’. Therefore in Very basic set theory the naming of individuals is logically first.
@@ 2. Logic and definiteness
In this section we prove logical formulas preserve definiteness. Formulas are denoted by $\Phi$, $\Psi$, $\ldots$ and are built up from formulas by means of negation $\neg\Phi$ (not $\Phi$) and conjunction $\Phi\wedge\Psi$ ($\Phi$ and $\Psi$). Negation takes precedence over conjunction. If $\Phi$ is a formula then $(\Phi)$ is a formula. The parenthesis (( and )) preclude ambiguity and improve legibility. Then $s\notin t$ is the formula $\neg(s\in t)$.
If $\Phi$ is a formula then $\forall s(\Phi)$ and $\exists s(\Phi)$ are formulas. $\forall s(\Phi)$ means ‘for every set it is the case $\Phi$’. The negation of $\forall s(\neg\Phi)$ is denoted by $\exists s(\Phi)$ thus $\exists s(\Phi)$ means ‘exists a set such that $\Phi$’. The parenthesis following $\forall s$ and $\exists s$ are mandatory. However, e.g. without ambiguity $\forall s\exists t(s\in t)$ is short for the formula $\forall s(\exists t(s\in t))$.
@@
$\Phi\vee\Psi$ ($\Phi$ or $\Psi$ or both) is short for the formula $\neg(\neg\Phi\wedge\neg\Psi)$,
$\Phi\rightarrow\Psi$ ($\Phi$ implies $\Psi$) is short for the formula $(\neg\Phi\vee\Psi)$,
@@
There are many rewrite rules: $\neg(\neg\Phi)$ is logically equivalent to $\Phi$, $\neg(\Phi\leftrightarrow\Psi)$ is logically equivalent to $(\Phi\leftrightarrow\neg\Psi)$ and $((\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi))$ is logically equivalent to $((\Phi\wedge\Psi)\wedge(\neg\Phi\vee\neg\Psi))$. More rewrite rules in (Mendelson [2015] p.15).
Definition.
Formula $\Phi$ is ‘definite’ if it is not the case that both $\Phi$ and $\neg\Phi$. $\Box$
Theorem.
The building up of formulas preserve definiteness.
Proof. For every $s$ and for every $t$ it is not the case that both $s\in t$ and $\neg(s\in t)$
therefore atomic facts and elementary propositions are definite.
If $\Phi$ is definite then it is not the case that both $\Phi$ and $\neg\Phi$. Thus it is not the
case that both $\neg(\neg\Phi)$ and $\neg\Phi$ therefore $\neg\Phi$ is definite.
If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise $(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ or
$(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $\exists s(\Phi\wedge\neg\Phi)$.
And $(\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi)$ is logically equivalent to
$((\Phi\wedge\neg\Phi)\wedge\Psi)\vee(\Phi\wedge(\Psi\wedge\neg\Psi))$ therefore if $\Phi$ and $\Psi$ are definite then
$(\Phi\wedge\Psi)$ is definite.
The logical connectives ‘or’, ‘implies’ and ‘if and only if’ are expressed in terms
of logical ‘not’ and logical ‘and’. Thus if $\Phi$ and $\Psi$ are definite then $(\Phi\vee\Psi)$,
$(\Phi\rightarrow\Psi)$ and $(\Phi\leftrightarrow\Psi)$ are definite.
If an atomic fact is analyzed as fully as possible constituents finally reached are
individuals. Atomic facts and elementary propositions are definite.
Therefore the building up of formulas preserve definiteness. $\Box$
3. Equality for sets
Experience shows a set ‘is equal to’ or ‘is not equal to’ another set. A set is determined by its constituents. Thus sets with different constituents are different sets, and the constituents of different sets are different. We define equality for sets.
@@ 3. Equality for sets
4. Axiom ‘specification’
  • if $u$ is free in $\Phi$ and $\Phi$ is definite then
Theorem.
If an individual is a constituent of set $s$ then exists set $v$ whose
constituents are all and only those individuals that are constituent of $s$.
Proof. Formula $(u\in s\wedge u\in u)$ is definite and is not the conjunction of mutually
exclusive properties. Apply specification to set $s$ to find set $v$
The formula applies to every $u$ thus we can choose $v$
$\exists v(v\in v\leftrightarrow(v\in s\wedge v\in v))$ therefore
Then $\exists v(v\notin v\vee v\in s)$.
Therefore $v$ is not an individual, or $v$ is an individual and a constituent of $s$. $\Box$
Theorem.
The set of all sets is nonexisting.
Proof. If $\exists u(u\in s\wedge u\notin u)$ then formula $(u\in s\wedge u\notin u)$ is definite and is not the
conjuction of mutually exclusive properties.
Apply specification to set $s$ to find set $v$
The formula applies to every $u$ thus we can choose $v$
Then $\exists v(v\notin v\wedge v\notin s)$.
Thus $v$ is not an individual and $v$ is not a constituent of $s$.
Therefore the set of all sets is nonexisting. $\Box$
@@ 5. Axiom ‘pairs’
Set $v$ is denoted by $\{s,t\}$, $s$ and $t$ are the only constituents of $\{s,t\}$.
Then $\{s,t\}=\{t,s\}$ thus $\{s,t\}$ is also called ‘an unordered pair’.
The ‘singleton’ $\{s\}$ of set $s$ is $\{s,s\}$. Then
@@
However, if $s\neq t$ and $(s\in t\wedge t\in s)$ then $s\notin s$ and $t\notin t$. Then exists set $\{s,t\}$ and the ‘vicious membership circle’ $(s\in t\wedge t\in s)$. That is contradictory to ‘the naming of individuals is logically first’. Therefore if $s\neq t$ it is not the case that
$\forall v(v\in\{s,t\}\rightarrow\exists u((u\in\{s,t\}\wedge u\in v)\wedge u\notin u))$.
The naming of individuals is logically first. Thus vicious membership circles do not exist. We prove ‘regularity’.
Theorem.
It is not the case that exists set $w$ such that
$\forall v(v\in w\rightarrow\exists u((u\in w\wedge u\in v)\wedge u\notin u))$.
Proof. A finite sequence of sets $s$, $t$, $u$, …, $z$ such that $s\in t$, $t\in u$, …, $z\in s$
is nonexisting otherwise there are vicious membership circles. And an ‘infinite’
sequence of sets $s$, $t$, $u$, $v$, … such that each term is a member of the previous
one $t\in s$, $u\in t$, $v\in u$ is nonexisting otherwise for every set $x$ belonging to
the sequence exists set $y$ belonging to that sequence such that $y\in x$.
That is contradictory to ‘the naming of individuals is logically first’. $\Box$
Formula $\neg\exists w\forall v(v\in w\rightarrow\exists u((u\in w\wedge u\in v)\wedge u\notin u))$ is definite. Therefore
  • (regularity) $\forall w\exists v(v\in w\wedge\forall u((u\in w\wedge u\in v)\rightarrow u\in u))$.
The formula applies to every $u$ thus we can choose $v$. Then the formula reads
$\forall w\exists v(v\in w\wedge((v\notin w\vee v\notin v)\vee v\in v))$, that is $\forall w\exists v(v\in w)$.
Therefore an ‘empty set’ is nonexisting. If an individual is a constituent of $w$ then $v$ is an individual. Otherwise $v$ is not an individual and $w$ and $v$ only have individuals in common.
@@ 6. Axiom ‘union’
Theorem.
Any sequence of different sets $s$, $\bigcup s$, $\bigcup(\bigcup s)$, … is finite.
Proof. Then $s\neq\bigcup s\rightarrow\exists t(t\in s\wedge t\notin t)$ otherwise $s=\bigcup s$.
@@
Therefore the sequence terminates. Thus exists set $w$ belonging to the sequence
such that $\bigcup w=w$. If $z=s\cup\bigcup s\cup\bigcup(\bigcup s)\cup\ldots$ $w$ then $z$ is a set and $\bigcup z=z$.
Set $z$ is called ‘the transitive closure’ of $s$. $\Box$
The definitions of equality and axiom pairs are only meaningful if exist sets the naming of which are logically first. Thus with individuals and regularity. And regularity is used with axiom union to prove that for every set the transitive closure exists. The proof needs no other axioms and no ‘set of all natural numbers’.